Find the Fahrenheit temperature of hot coffee at 75oC by solving the equation 75 =5/9(f - 32)for F
Answer by checkley77(12844)

75 =5/9(f - 32)
75=5f/9+5*-32/9
75=5f/9-160/9
5f/9=75+160/9
5f/9=75+17.78
5f/9=92.78
5f=9*92.78
5f=835.03
f=835.03/5
f=167 degrees is the answer.